### v1

# MpoxMamba: A Grouped Mamba-based Lightweight Hybrid Network for Mpox Detection

Abstract
Due to the lack of effective mpox detection tools, the mpox virus continues to spread worldwide and has once again been declared a public health emergency of international concern by the World Health Organization. Deep learning-based mpox detection tools are crucial to alleviate mpox outbreak. However, existing methods have difficulty in achieving a good trade-off between detection performance, parameter size, and model complexity, which is crucial for practical applications and widespread deployment, especially in resource-limited scenarios. Given that the success of Mamba in modeling long-range dependencies and its linear complexity, we proposed a lightweight hybrid architecture called MpoxMamba. MpoxMamba utilizes deep separable convolutions to extract local feature representations in mpox skin lesions, and greatly enhances the model’s ability to model the global contextual information by grouped Mamba modules. Experimental results on two widely recognized mpox datasets demonstrate that MpoxMamba outperforms existing mpox detection methods and state-of-the-art lightweight models. We also developed a web-based online application to provide free mpox detection services to the public in the epidemic areas (http://5227i971s5.goho.co:30290/). The source codes of MpoxMamba are available at https://github.com/YubiaoYue/MpoxMamba.
Index Terms:
Mpox Detection, Deep Leraning, Vision Mamba, Lightweight Model.I Introduction
The fast-spreading mpox (monkeypox) poses a huge threat to global health [1, 2]. Over 120 countries have reported mpox between Jan 2022 and Aug 2024, with over 100,000 laboratory-confirmed cases reported and over 220 deaths among confirmed cases. Given the upsurge of mpox in the Democratic Republic of the Congo (DRC) and a growing number of countries in Africa, WHO declares mpox outbreak a public health emergency of international concern on August 14, 2024 [3], highlighting the urgent need for rapid and accurate mpox detection tools.
Although accurate, traditional mpox detection methods, such as Polymerase Chain Reaction testing, are often limited by equipment, cost, and time efficiency, which is not conducive to early detection and large-scale screening of mpox, especially in areas with limited medical resources [4]. In recent years, deep learning technology has made significant progress in the field of medical image processing. Deep learning systems based on skin lesion images have been proven to be a reliable solution for rapid and accurate mpox detection. Although most existing methods have successfully utilized convolutional neural networks (CNNs) to detect mpox [5, 6], the inherent locality of the convolution operator limits the model’s ability to capture long-range dependencies in skin lesion images. Long-range dependencies are a prerequisite for the formation of contextual information, which is crucial for accurate skin disease classification [7]. Meanwhile, due to the intra-class variability and inter-class similarity in skin disease images, an effective mpox detection model should effectively integrate local features and global contextual information[8, 9]. To address this problem, some studies have used the Vision Transformer (ViT) to enhance the CNN-based model’s ability to model long-distance dependencies [10, 11]. Unfortunately, ViT’s detection performance relies on a large amount of pre-training data and the quadratic complexity of the self-attention mechanism results in expensive computational costs, which hinders the real-time application of mpox detection in large-scale screening scenarios. More importantly, existing high-accuracy mpox detection models often have large parameter sizes and are not suitable for widespread deployment, especially on devices with limited computing and memory resources (Such as smartphones and development boards).
Recent studies have shown that Mamba with linear computational complexity exhibits excellent performance in modeling long-distance dependencies[12]. Inspired by this, we hypothesize that Mamba can enhance lightweight CNN’s ability to extract global contextual information, thereby promoting the practical application of accurate and efficient mpox detection tools, which has important realistic meaning for epidemic prevention and control and global public health security. In this work, we investigated the potential of lightweight hybrid networks based on Mamba-CNN for mpox detection, aiming to achieve a good trade-off between detection performance, parameter size, and model complexity. To the best of our knowledge, this study represents the first attempt to use Mamba for intelligent mpox detection.
The contributions of this work can be summarized as follows: (1) We first proposed a lightweight hybrid network based on Mamba and CNN for efficient mpox detection, named MpoxMamba. The parameter size and model complexity (FLOPs) of MpoxMamba are only 0.77M and 0.53G. (2) We introduced a novel Grouped Mamba-based Local-Global Feature Fusion Block (GMLGFF). GMLGFF uses the grouped Mamba to enhance the model’s ability to capture the long-range dependencies, helping the model perceive and aggregate scattered local features and global contextual information in mpox lesions effectively. By splitting feature maps and parallel processing, GMLGFF greatly reduces model complexity and parameter size. (3) Experimental results on two widely recognized mpox datasets show that the MpoxMamba surpasses the state-of-the-art (SOTA) lightweight networks and existing methods. (4) A web-based online application was developed to provide free mpox detection services to the public in the epidemic area.
II The Proposed Method
II-A The Overall architecture of MpoxMamba
Fig. 1a shows the overall framework of MpoxMamba. We first used a convolution layer and inverted residual (InRes) block [14] (Fig. 1b) to capture low-level features such as the appearance, shape, and color of the rash image, resulting in a 32×112×112 feature map. Then, an IRB block with stride=2 is used to quickly reduce the spatial dimension and optimize the extracted features. Stage 2 and Stage 3, constructed by stacked GMLGFF Blocks, are used to comprehensively extract local features and global context information. The InRes Blocks (stride=2) (Fig. 1c) are used to downsample the outputs of Stage2 and Stage3, outputting two feature maps with resolutions of 128×28×28 and 256×14×14, respectively. A point-wise convolution is used at the end of the network to map the feature map to a high-dimensional space, enriching the feature representations. Finally, the adaptive average pooling and linear layers are used to detect mpox and non-mpox. The activation function and input size of MpoxMamba are set to SiLU and 224×224×3, respectively.
II-B Vision Mamba Layer
The Vision Mamba (VM) layer integrates a cross-scan module [15] to assist the model in traversing the spatial domain of image feature maps from four corners all across the feature map to the opposite location, which ensures that each pixel in a feature map integrates information from all other locations in different directions, resulting in a global receptive field without increasing the linear computational complexity. The S6 Block[13] is used to process sequences from all directions to ensure that the model can effectively model long-distance dependencies. Finally, the output features from the four directions are merged through a scan-merging operation to construct the final 2-D feature map, resulting in a final output of the same size as the input. Fig. 1e shows the modeling mechanism of the VM layer. The pseudo-code for the S6 block is presented in Algorithm 1.
II-C Grouped Mamba-based Local-Global Feature Fusion
The proposed GMLGFF block (Fig. 1d) first employs depthwise separable convolution (local representation module) to extract local feature representation without changing the size of the feature map. Then, the output of the local representation module () is evenly divided into four sub-feature maps (), and four Vison Mamba layers are used to model the sub-feature maps to extract global feature representation (global representation module). In the feature fusion module, Efficient Channel Attention [16] is used to adaptively adjust the output of the local representation module, thereby highlighting the key local features of mpox lesions while suppressing irrelevant information. Subsequently, the optimized local features are concatenated with the output of the global representation module along the channel dimension (), and the concatenate feature map is modeled using point-wise convolution, which ensures that the model can consider both local details and global context information and further improves the interaction between local-global feature representations. A shortcut connection is used to keep the original features, thereby reducing the risk of information loss in deeper layers.
III Experiments and Results
III-A Dataset and Evaluation Metrics
Two widely recognized mpox datasets are used in this work: Monkeypox Skin Lesion Dataset (MSLD) [17, 18] and Monkeypox Skin Images Dataset (MSID)[5]. MSLD is a binary classification dataset containing 102 mpox images and 126 non-mpox images. MSID is a multi-class dataset containing 279 mpox images, 91 measles images, 107 chickenpox images, and 293 normal skin images. Here, based on the characteristics of medical images, we used overall accuracy (OA), sensitivity, and specificity as evaluation metrics.
III-B Implementation Details
We did not apply any pre-training strategy or data augmentation during the experiment to demonstrate as much as possible that the model performance only benefits from the Mamba-CNN-based model. Five-fold cross-validation was used to fully evaluate the model’s performance and generalization ability, reducing the risk of overfitting. We used the Ubuntu 22.04 operating system with NVIDIA GeForce RTX 3070 and Pytorch to complete the entire experiment. During the training process, we set the epoch and batch size to 100 and 16 respectively, and used the AdamW optimizer (learning rate=0.0001, 1=0.9, 2=0.999 and weight decay=1e-4) and cross-entropy loss to optimize the model. To reflect the superiority of the MpoxMamba, some SOTA lightweight networks and existing methods were selected as reference models. To ensure fairness, the network architecture was the only different factor in the model comparison experiment.
III-C Comparison of Results
Table LABEL:table1 reports the results of various metrics of MpoxMamba and the reference models. Since MSLD is a binary classification dataset, we only calculate the sensitivity (abbreviated as Se) and specificity (abbreviated as Sp) of the model in detecting mpox. For MSID, we calculate the average sensitivity and average specificity of all categories.
The results show that the OA of MpoxMamba reached 82.47% (MSLD) and 81.71% (MSID), respectively, reaching SOTA performance among all models. Compared with MonkeyNet, the OA of MpoxMamba increased by 2.17% and 0.66%, respectively. Except for Se (MSLD), MpoxMamba surpassed MonkeyNet in other metric results. It is worth noting that the parameter size and FLOPs of MpoxMamba are only 0.77M and 0.53G, respectively, which are 95.74% and 87.70% lower than MonkeyNet, respectively. Compared with the FasterNet-T1, MpoxMamba improves OA by 7.47% and 8.35% on MSLD and MSID, respectively, achieving a great performance improvement while maintaining lightweight. A comprehensive analysis of the results in Table 1 demonstrates that MpoxMamba has shown great competitiveness in various metrics and achieved the best trade-off between detection performance, parameter size and model complexity.
Model | Paras | FLOPs | OA(MSLD) | Se(MSLD) | Sp(MSLD) | OA(MSID) | Se(MSID) | Sp(MSID) |
Mobileone-S0*[19] | 4.27M | 1.07G | 73.58% | 55.46% | 88.12% | 70.01% | 57.45% | 88.74% |
MobileViT-XS*[20] | 4.94M | 1.43G | 74.92% | 70.36% | 78.56% | 72.45% | 60.92% | 89.94% |
EfficientFormerV2-S1*[21] | 5.74M | 0.66G | 71.92% | 60.90% | 80.90% | 63.52% | 51.85% | 86.48% |
FastViT-t8*[22] | 3.26M | 0.53G | 75.00% | 66.64% | 81.76% | 73.36% | 63.07% | 90.19% |
FasterNet-T1*[23] | 6.32M | 0.85G | 77.19% | 69.54% | 83.36% | 73.23% | 62.34% | 90.11% |
MonkeyNet[5] | 18.10M | 4.31G | 79.44% | 77.36% | 81.04% | 81.05% | 72.00% | 93.02% |
PoxNet22[6] | 24.35M | 5.70G | 74.50% | 67.64% | 80.1% | 72.09% | 62.44% | 89.89% |
MpoxNet[24] | 8.86M | 1.37G | 74.94% | 67.46% | 80.96% | 76.51% | 67.35% | 91.28% |
ResNet50[25] | 23.52M | 4.10G | 76.25% | 74.18% | 77.82% | 74.82% | 66.34% | 90.85% |
SwinTransformer-T[26] | 27.52M | 4.49G | 77.63% | 70.64% | 83.32% | 75.32% | 67.48% | 91.33% |
ShuffleNetV2-X20 [27] | 5.35M | 0.58G | 72.33% | 50.64% | 89.76% | 69.89% | 59.70% | 88.83% |
EfficientNet-B1[28] | 6.52M | 0.58G | 71.02% | 61.46% | 78.64% | 68.79% | 56.36% | 88.37% |
MpoxMamba(Ours) | 0.77M | 0.53G | 82.47% | 75.46% | 88.12% | 81.71% | 74.07% | 93.37% |
III-D Ablation Experiment
We used MSLD to conduct ablation experiment. The entire ablation experiment started with the base model (the base model means that we removed the global representation module and the feature fusion module in GMLGFF Block). We analyzed the impact of the number of branches of Grouped Mamba on model performance, parameter size, and GFLOPs. Table II shows the ablation experiment results in detail. It can be seen from the results that the VM and Fusion modules significantly improve the OA of the basic model but inevitably increase the model complexity and parameter size. Fortunately, increasing the number of branches of Grouped Mamba improves the model accuracy, parameter size, and model FLOPs.
Model | VM | Fusion | VM=2 | VM=3 | VM=4 | OA | Paras | GFLOPs |
Basic Model | ✗ | ✗ | ✗ | ✗ | ✗ | 72.18% | 0.29G | 0.34M |
Model2 | ✓ | ✗ | ✗ | ✗ | ✗ | 77.63% | 0.69G | 0.96M |
Model3 | ✓ | ✓ | ✗ | ✗ | ✗ | 80.39% | 0.82G | 1.07M |
Model4 | ✓ | ✓ | ✓ | ✗ | ✗ | 82.03% | 0.63G | 0.87M |
Model5 | ✓ | ✓ | ✓ | ✓ | ✗ | 81.72% | 0.56G | 0.81M |
MpoxMamba | ✓ | ✓ | ✓ | ✓ | ✓ | 82.47% | 0.53G | 0.77M |
III-E A Web-based Mpox Detection Application
To promote the practical application of deep learning technology, we utilized the MpoxMamba to develop a web-based online application (Fig. 2). Users only need to upload a pre-taken rash image, and they can obtain real-time diagnosis results through the Internet. In order to improve the security of the application, we also used Grad-CAM [29] to visually explain the model’s decision-making results. In addition, medical knowledge and medical advice on various diseases and reference images are displayed on the result page to raise public health awareness. We hope that this work can provide the best possible help to the epidemic areas.
IV Conclusion
In summary, this work proposed a lightweight Mamba-CNN hybrid model that can efficiently detect mpox lesions in rash images. Compared with existing methods, MpoxMamba achieves SOTA performance while maintaining the lowest parameter size and model complexity. Future research will focus on exploring effective data augmentation techniques and pre-training strategies (such as self-supervised learning and transfer learning) to further improve the performance of MpoxMamba.